This is not a NASA Website. You might learn something. It's YOUR space agency. Get involved. Take it back. Make it work - for YOU.

Shuttle Mir Science Program Status Report 3-9 April 1997

By Keith Cowing
April 9, 1997
Filed under

Shuttle Mir Science Program / NASA 4 / Mir 23 Status
Date: April 3, 1997 – April 9, 1997
Mission Days: 82-88 (NASA 4) / 53-59 (Mir 23)
Mir Systems
Telemetry – The OMOTsAO unit failed yesterday; this device stores uplinked executable commands, in particular, the command for the MIPS downlink. That means that there will be no MIPS downlink available until the system is repaired. Repair activities on this unit are scheduled for 14 April.Suggestion will be made to Jerry to use the Packet system. (Jim Medford and Ed Walters (MOD) are trying to understand what the MOTsA is. The SM IPT is going to see if a similar unit is in phase 2. This is the first appearance of this in our documentation.
Progress – Successful Auto-docking of the Progress occurred last night at 20:30 DMT. The crew was directed to open the hatch early last night and to temporarily suspend using candles, utilizing the O2 from the Progress. Not sure if ethylene glycol was carried on this flight or how much in case they need to replace the amount leaked out from the thermal system. They were not sure if another solid oxygen generator unit was flown up. The POSA will get the manifest and send it to me in the next week or so.
I have asked for a contact on understanding how they use the O2 from the Progress to replemish the MIR oxygen. Jim Medford is the contact to get this info. The Russians apparently use this process all the time (ie, not just as a result of the recent failures). I therefor suspect we will do it for Station and suggest it would be helpful to understand this capability. I believe we have maintained a record of the atm content pre and post Progress docking.
Repair plan – Two crew members working on removing the leaking heat exchanger from the VGK loop in Kvant 1. The crew will cut and cap the pipes leading in and out of the leaking heat exchanger in the Kvant 1.After the loop has been leak checked, the Vozdukh system will be reactivated. The goal is to complete this activity and reactivate the loop by the end of the 10th. The ground personnel are confident that the remaining three heat exchangers will be adequate to handle the system load.
The other crew member is working on the Elektron system in Kvant 2. The original goal was to repair and re-start the unit on the 9th. We do not have a status on this yet.

No word yet if IFM has actually begun. Will get status on Friday. Not sure which of the inflight maintenance activities Jerry will be involved with. We should note this activty for station as we work toward an INTEGRATED crew. Note that Jerry is having to give up US science for MIR systems repair. Also note later in this summary that Jerry’s EVA scheduled for the 29 is in jeopardy of cancellation. This is a vehicle risk mitigation experiment (RME).
Leaks – Leak isolation and detection on the KOB 1 & 2 loops in the Core Module will take 4 days. The preliminary plan is for this to occur from the 11th through the 14th, although the 12th is a major holiday (Cosmonautics Day), so the schedule is likely to change. A total of 11 leaks have been identified in the KOB -1 and -2 loops in the base block, and the heat exchanger for the VGK in Kvant. None of these loops are operational. In their search for the leak, the crew reported that Othe quality of many of the pipes is not goodO. The ground estimates that a total of 3.8 liters has leaked out this mission.
This is the largest leak they have seen in space. There was a leak in December 1995 of about 1.5 ltr. Need to get the Doc’s and toxic folks in on this. Apparently the amount of ethylene glycol combined with the hight temperatures (see below) and its location to fans (aerosolyze?) worsen the effects. Studies to date did not take the last two functions into account.
Also, floresent dye was flown up to check the leaks by next week.
Temperatures – The Kvant 2 thermal loop repair and check-out has been completed and the loop has been reported as up and operating. The temperature in Kvant 2 is reported to be 20 – 21.5 deg C, warmer in the area near the gyrodines (approx 26 deg C). Spektr showed a max of 27 deg C, and Kristall (where Jerry is working out on the treadmill) 26 deg C. As a result of redirection of airflow into the base block and Kvant, the temperature in the base block is now 32, and in Kvant 25 deg C.
The Core Module temp reached as high as 40 deg C, until the crew redirected some internal ducting. The temperature is now back in the low 30Os.
Science Status
The Human Life Sciences Humoral Immunity (HI) investigation concluded on April 4 with a blood and saliva collection. This experiment assesses the production and action of antibodies in response to an antigen vaccination during spaceflight. Blood and saliva samples are collected at timed intervals to measure antibody production and determine the effectiveness, extent and time course of the antibody response.
The NASA 4 crewmember was scheduled to process two sample pairs for the Microgravity Queens University Experiment in Liquid Diffusion (QUELD). This experiment evaluates the diffusion coefficients of select metals in microgravity, and uses the Microgravity Isolation Mount (MIM) as an isolation platform from the Mir vibrational environment.
The Tissue Equivalent Proportional Counter (TEPC) readings were communicated to the ground on April 3. TEPC measures radiation in the living space of the Mir station. It is limited to measuring the general area where it is located onboard, in the Spektr module. It measures radiation cumulatively, but also has a real-time Liquid Crystal Display readout.
The weekly Kanas “Interactions” questionnaire, which profiles mood states and interpersonal group environments, was completed as scheduled on April 9. This activity is scheduled for both the NASA and Mir crews on Wednesdays throughout NASA 4.
The NASA 4 crewmember initiated his second block of the Sleep Investigations on March 31. This block will continue for twelve nights.On the fourth and fifth day of this block, April 3rd and 4th, the NASA 4 crewmember will take blood samples. This investigation studies the effects of microgravity on sleep/wake cycles, vestibular adaptation, and the immune system.
A Microbiological Assessment of the Mir Station was performed on April 2.Samples of the air and Mir surfaces. Follow-up assessments of these cultures will be performed on April 4 and April 7. This investigation studies the microbiological life on Mir station
Processing of the Thermoluminescent Dosimeters (TLDs) for the Effective Dose Measurement at EVA experiment was scheduled for April 8. This experiment utilizes TLDs to characterize the Mir radiation environment. The TLDs will be continue to be processed at regularly scheduled intervals.
A Space Acceleration Measurement System (SAMS) session was scheduled to record the accelerations on the Mir space station during the Progress 234 docking on April 8.
A Mir Structural Dynamics Experiment (MiSDE) session was also scheduled for the docking of Progress 234 on April 8. The MiSDE experiment measures accelerations on Mir during certain dynamic events such as vehicle docking, crew exercise, and thruster firing.
The NASA 4 crewmember was planned to evaluate the Gas Analysis System (GASMAP) hardware on April 9. This hardware is checked out approximately every 30 days during the NASA/Mir Program and is used to characterize physiological parameters during exercise.
SIGB – Jerry reported after performing the SIGB diagnostic radiogram, the SIGB lights came on. He did not know what was different from the first time he tried. He stated the SIGB is fully functional and ready for NASA 5 experiments.
Orientation – The crew was not able to get the software to load for this experiment. Clarification came from Susan Minor that the PCMCIA cards in use for this experiment are of the type currently used by the MIPS system and have shown resistance to radiation. The Russian curator has been in contact with the French PEDs on this experiment. At the current time, no further Orientation sessions or troubleshooting will be scheduled for the remainder of NASA 4.
GASMAP – A radiogram was uplinked this week which will have Jerry make permanent changes to the contingency procedures in his FDF. He will perform a partial GASMAP checkout and if required use the new procedures to determine whether or not a vacuum problem has occurred with the unit. If so, a radiogram with corrective procedures to the FDF Roughing procedures will be uplinked and Jerry will be asked to move the unit to the Krystall module to access the vacuum port to perform the roughing procedures.
MiSDE – The thruster firing and Progress docking radiogram was uplinked.
GASMAP – The GASMAP radiogram was submitted to the Russian procedure specialists.
Operational Items Jerry reported that he would prefer to continue and finish QUELD, vice starting up a new experiment. The current plan is for Jerry to assist with system repairs in the near future, utilizing the limited remaining time to finish QUELD. EVA prep and packing constrain the remaining time available.
Victor Blagov informed us that the Energia management has set a deadline of the end of April to successfully complete the maintenance activities. If these activities are not finished, then the EVA may be in jeopardy.Furthermore, he requested that we clear Jerry’s calendar of science through Friday, 11 April, in order to allow him to work on the maintenance activity. As a result, Jerry will perform at only the Sleep experiment and CSST on Friday.
Jerry on a later pass said that he does not have any problem performing system maintenance activities, but wanted to make sure that the cancellation/delay of science is analyzed carefully.
Stowage/Transfer Operations
Nothing to report.
All three Mir crewmembers performed MK-6, Body Mass Measurement, MK-7, Calf Volume Measurement, MK-27 Urinalysis, and portions of MK-12, Reflotron blood analysis. The MK-12 results indicate a compromise of the integrity of the Reflotron kit, probably due to elevated temperatures in the core module. The crew may be asked in the near future to repeat the test.
Due to the current Mir environmental conditions, the crew has been monitoring pulse rate, blood pressure, skin temperature, and blood oxygen saturation twice per day and relaying the results to the ground.Crewmembers have also been instructed to discontinue their nominal exercise regime. A modified exercise schedule of walking on the treadmill and using the expanders in the Kristall is allowed. Recommendations on how to prevent heat related physical problems were also relayed to the crew.
Jerry has been receiving all toxicology information regarding ethylene glycol as related to the onboard situation. On 6 April he requested to speak to the NASA crew surgeon regarding some of the information sent onboard. No issues resulted.
The Crew Status and Support Tracker (CSST) was completed on 4 April. The file has not been received since a MIPS downlink has not been received since the completion of this activity.
T + 2 and T + 5 colony counts were performed for Microbial Air Samples and Surface Samples.
Jerry participated in an audio only non-private comm with the flight surgeon on 6 April, and with the Ops Lead and flight surgeon on April 9th. The events included approximately 20 minutes and 5 minutes, respectively, of good comm.
PAO Status
Last week Jerry was interviewed by CNN International in Spain. The event was aired on NASA TV. CNN International also aired the interview live, while CNN-domestic taped it for rebroadcast during the day. Due to problems with MCC-M receiving the TV signal from the ground stations, TV coverage was sporadic and of poor quality.
Following this, the decision was made to not attempt live TV for future PAO events until the satellite is available.
We have arranged for a visit by Craig Covault, reporter for Aviation Week and Space Technology, who plans to spend two days in the TsUP observing the MOST team at work and interviewing various Russian officials. Covault will be in Moscow for two weeks, working on a story about the International Space Station.
The Moscow bureaus of CBS News, CNN, Reuters, WTN and the Christian Science Monitor covered the docking of Progress with Mir from the TsUP. Tony Sang was interviewed by Reuters, WTN and the Christian Science Monitor.
A PAO event is planned for April 11 from 1600-1618 DMT. It will be a press conference with reporters at KSC, JSC and HQ. Jerry has been alerted to the fact that April 12 is Cosmonautics Day (the anniversary of Yuri Gagarin’s 1961 flight), and it also is the anniversary of the first shuttle flight in 1981. He also is aware of the “feeding frenzy” by the U.S. news media over recent problems on shuttle and Mir.
PAO Planning
Next week’s event with Jerry is tentatively planned for Thursday, April 17.
Client and time are TBD.

Date                           Discipline      Activity
10 April                               HLS             Sleep Block 2, Day 11
11 April                               HLS             Sleep Block 2, Day 12
                                SMP             CSST Questionnaire
                                               MIPS Prep & Downlink
12 April                               HLS             Sleep Block 2, Day 13
                                MG              BTS Filter cleaning
13 April                               PFC             Private Family
14 April                               ISS             MiSDE IVA
                        MIPS Prep. & Downlink
                               MG              QUELD
15 April                               MG              DCAM Photography
16 April                               HLS             Interactions
                               SMP             Formaldehyde Monitoring
                                               MIPS Prep. & Downlink 

Mission Science Office
(713) 483-1085 or (713) 280-2755
SMSP Operations
JSC Payload Operations Support Area
(713) 483 – 1224

NASA Watch founder, Explorers Club Fellow, ex-NASA, Away Teams, Journalist, Space & Astrobiology, Lapsed climber.